# Automated MRI Tumor Segmentation using hybrid U-Net with Transformer and Efficient Attention

### v1

@@ 1 INTRODUCTION
Improper classification and identification of brain tumors is a major cause of high mortality rate worldwide in afflicted patients. An estimated 4,407 deaths from brain and central nervous system cancers occurred in Pakistan in 2022, highlighting the critical need for improved diagnosis and treatment methods [4]. Magnetic resonance imaging (MRI) offers superior soft-tissue contrast, making it the modality of choice for identifying brain tumors and lesions [Essig2012MR]. Precise segmentation of these abnormalities is vital for both early detection and the formulation of appropriate treatment strategies. An ischemic stroke results from the obstruction of a cerebral blood vessel, which deprives brain tissue of essential oxygen and nutrients. Such an event represents a critical medical emergency that may lead to irreversible neurological damage or fatality[Adams1993Classification]. Manual segmentation remains the clinical gold standard for delineating tumor boundaries, despite being time-consuming and subject to inter-observer variability[Warfield2004STAPLE]. Medical experts in local hospitals still use manual segmentation of tumor regions to mark areas and boundaries of damaged brain tissue.
@@
Manual segmentation has its own daunting challenges which impact the patient directly. It is a time-consuming process and often a difficult task left out for the oncologists. Unlike well-defined organs (eyes, brainstem), tumors are irregular and unpredictable. They have inconsistent shape, size, boundaries and often blend with surrounding tissues, making detection harder. Dynamic growth and irregular borders complicate precise segmentation, and the contouring is often subjective to the view of the oncologist. There is a need for automated methods for tumor segmentation that can offer a more efficient and automated approach to diagnosing and treating tumors[Litjens2017Survey]. Supervised deep learning methods like CNNs have contributed to image segmentation tasks in medical physics greatly over the years, especially reflecting in the BRATS[menze2015multimodal] challenges annually. UNet [unet] is one of the most used deep learning architecture in medical image segmentation. The UNET architecture consists of an encoder and decoder block linked with a bottleneck layer in a U shape (hence the name UNet). The purpose of the encoder block is to downsample the input image using convolutional and pooling layers to condense the information in it by contracting the spatial resolution and depth structures, passing it to a bottleneck layer that connects it to the decoder layer. The decoder layer then upsamples the condensed image to an output label mask using Transpose Convolutions.
@@
While pure CNNs excel at capturing local features, they struggle with long-range dependencies. A recent survey of self-supervised learning mechanisms for Vision Transformers demonstrates how SSL-pretrained ViTs can greatly reduce labeled data requirements while preserving downstream segmentation accuracy [survey_ssl_vit_2408.17059]. At the same time, an extensive review of Vision Transformer–based models for medical image segmentation reports that hybrid CNN–ViT architectures consistently outperform pure CNNs on MRI and CT benchmarks by better modeling global context [survey_medseg_vit_2312.00634].
@@
While models like UNET and its variants have shown remarkable success in the BRATS challenges, they face significant hurdles when applied to real-world clinical data, particularly in resource-limited settings. The reliance on curated datasets such as BRATS poses critical challenges for models aimed to be deployed locally in hospitals for radiation oncology treatment planning. Public datasets often fail to capture the heterogeneity of local patient populations, imaging protocols, and scanner variations. In addition, Models trained on BRATS or similar datasets may not generalize well to localized datasets, which are smaller in size and exhibit greater variability[Guan2022Domain]. Data privacy concerns and regulations restrict the availability of medical data, thus limiting the prospect of such studies to test out models on vast amount of data[GDPR2016].
@@ 2.1 Dataset and Data Preparation
The dataset used in this study was directly collected from a local hospital, offering a unique opportunity to evaluate our segmentation method on authentic clinical data. Unlike standardized benchmarks such as the BraTS dataset, which is employed in state‐of‐the‐art approaches like nnU-Net [dataset1] and UNETR [dataset2] or the MICCAI LiTS [bilic2019liver] dataset used in H-DenseUNet [HDENSENET], our dataset reflects the natural variability encountered in routine clinical practice. The inherent differences in imaging protocols and patient demographics introduce challenges for segmentation models, while simultaneously enhancing the real-world relevance of our findings.
As summarized in Table 1, the dataset comprises MRI scans from six patients, with four cases acquired using T1-weighted sequences and two with T2-weighted sequences. Each patient contributed approximately 175 slices, yielding a total of 700 T1-weighted and 350 T2-weighted images. All images were acquired at a voxel resolution of acquired at a voxel resolution of . The raw data, stored as DICOM files, were initially converted into volumetric images using the open-source PyDICOM library [pydicom]. Expert-provided segmentation masks were then transformed into binary label maps, and the volumetric data were sliced into 2D images, each paired with its corresponding mask. These processed images were subsequently organized into TensorFlow data pipelines to facilitate efficient model training and evaluation.
@@
Given the modest size of the original dataset (1̃,000 images), an extensive data augmentation strategy was implemented to enhance both data diversity and model robustness. As illustrated in Figure 2, several data augmentation techniques including flips, rotations, Gaussian blur, and contrast shifting are applied that expand the size of the dataset to 6,080 images. Such augmentation methods have been shown to improve the generalization and accuracy of deep learning models in medical imaging tasks [dataset4].
@@ 2.2 Proposed Methodology
Vaswani et al [attention] introduced the transformer architecture, which employs a self attention mechanism to process sequences concurrently, yielding high efficiency. This approach enables the network to attend to the most pertinent regions of an image by mapping a query against collections of keys and values to produce an output. A compatibility function between the query and each key computes weights for the associated values, which are then aggregated through a weighted sum to form the final output [attention]. By assigning attention weights in this way, the model captures long range dependencies and attains a wider receptive field than convolutional neural networks, thereby integrating broader contextual information. Combining the merits of both transformer and convolutional designs, hybrid models have shown remarkable performance improvement over baseline models.
Hybrid architectures have emerged as a significant advancement in deep learning, particularly for medical image segmentation tasks[chen2021transunet]. By leveraging the strengths of multiple methodologies, these architectures aim to improve performance and address inherent limitations in conventional models. Traditional UNet architectures, renowned for their inductive bias and efficient handling of small datasets, often struggle to capture long-range dependencies[unet]. Subsequently, transformer-based models excel in modeling global relationships but demand extensive annotated data, making them impractical for resource-constrained environments[dosovitskiy2020image]. Hybrid models offer a balanced solution by integrating the complementary features of these approaches[dataset2].
Several hybrid architectures have been proposed in the literature. The Hybrid UNet Transformer (HUT) [1] combines convolutional layers with transformers, introducing transformers at skip connections to enhance receptive fields. While effective in improving attention mechanisms, its reliance on large annotated datasets limits its applicability to resource-constrained settings. Similarly, the MaxViT-UNet framework [3] integrates MaxViT blocks into an encoder-decoder structure, achieving significant parameter efficiency and local-global feature modeling. However, the computational needs and resources required for such models make them less suitable for deployment in under-resourced environments. Another notable approach, the Two-Track UNet [HTTU], utilizes dual convolutional tracks to address class imbalances, yet its reliance on specific kernel configurations reduces adaptability to heterogeneous datasets. Finally, H-DenseUNet [HDENSENET] fuses 2D and 3D DenseUNets to handle intra- and inter-slice features, achieving strong performance on liver and tumor segmentation tasks. Despite its effectiveness, the model’s dependence on large-scale public datasets and computational resources creates challenges for real-world clinical deployment.
@@
Our proposed hybrid architecture (see Figure 4) builds upon the canonical U-Net framework to achieve precise segmentation of clinical images while addressing the limitations of resource-constrained training environments. In the Encoder, we utilize a ResNet-50[resnet50] backbone from the [yakubovskiy2019segmentation] library, pre-trained on ImageNet[imagenet]. This design choice is motivated not only by ResNet-50’s[resnet50] proven ability to extract robust hierarchical features from input images but also by the practical necessity to reduce training time and computational cost on Kaggle GPUs, which impose strict limits on available resources. The encoder produces feature maps at four distinct stages: Stage 1 yields a tensor capturing low-level details (edges and textures), Stage 2 outputs a tensor representing medium-level features, Stage 3 produces a tensor encapsulating high-level information, and Stage 4 delivers a tensor that abstracts semantic content. To further enhance these representations, each encoder block is augmented with Squeeze-and-Excitation (SE)[hu2018squeeze] and Convolutional Block Attention Modules (CBAM)[woo2018cbam]. The SE block, for example, recalibrates the input by first computing a channel descriptor via global average pooling, then passing it through two FC layers with a reduction ratio r (typically 16): and finally re-weighting the input by performing element-wise multiplication . This mechanism enhances the network’s sensitivity to clinically important features, such as subtle tumor boundaries.
@@
At the network’s core, the Transformer Bottleneck is employed to capture long-range dependencies that conventional convolutional layers cannot. Here, the deepest encoder output is a tensor, that is reshaped into a sequence of 1024 tokens (i.e., ) and processed through a series of four transformer blocks. Each block applies multi-head self-attention[attention] (as described in the previously provided Equation 1 and a feedforward network with 1024 hidden units, along with dropout (0.1) and layer normalization. This design significantly broadens the receptive field and integrates global contextual information, which is essential for delineating diffuse and irregular tumor regions.
@@
In the Decoder, the network gradually upsamples the feature maps to restore the original resolution. At every upsampling stage, skip connections fuse high-resolution encoder features into the decoder output. Before fusion, the skip connection features are refined using CBAM to recalibrate channel and spatial responses. Moreover, an efficient attention[shen2021efficient] mechanism is applied within the decoder to dynamically emphasize the most informative features. In this module, the input tensor is first reduced in channel dimension using a convolution, followed by an additive branch and a ReLU activation; a subsequent convolution, followed by a sigmoid activation, yields attention coefficients that modulate the original feature map as , thereby filtering out irrelevant information. To further refine these features, each decoder block incorporates additional SE blocks and ResNeXt[resNeXt] modules. The ResNeXt block leverages grouped convolutions to capture multi-scale spatial patterns efficiently. It first reduces channel dimensions via a convolution, then processes spatial information using a depthwise convolution, and finally expands the channel dimension with another convolution. A parallel shortcut branch ensures that the block benefits from residual learning, and a final ReLU activation completes the operation.
@@
In the Decoder (right, shown in green), the network gradually upsamples the feature maps back into an image of original resolution. Skip connections, as indicated by the dashed arrows labeled “Attn Gate” in Figure Figure 4, fuse high-resolution encoder features with the decoder output. Prior to fusion, each skip connection is refined using CBAM, and then the decoder applies an efficient attention mechanism to selectively emphasize the most informative features. Furthermore, the decoder interleaves additional SE blocks and ResNeXt blocks; the latter leverages grouped convolutions[resNeXt] to capture multi-scale patterns efficiently, ensuring that both local details and global context are preserved. This comprehensive integration of modules enables our architecture to balance the data efficiency of convolutional networks[unet] with the long-range image dependency modeling of transformers[chen2021transunet], thus achieving robust segmentation performance even on limited clinical datasets.
@@ 2.3 The Training of Hybrid Architecture
where is a hyperparameter empirically optimized to balance the contributions of the two losses. This hybrid formulation has demonstrated significant effectiveness in balancing global context with fine-grained segmentation accuracy in medical image segmentation tasks [Milletari].
@@ 2.3.1 Evaluation Metrics
The segmentation performance is assessed based on some evaluation metrics, we employed several metrics commonly used in medical imaging studies [dataset1, HTTU, HDENSENET]. These metrics include Dice Score, which is a measure of the degree of overlap between predicted and ground truth image, similar to Intersection over Union(IoU), which quantifies the ratio of the intersection to the union of predicted and actual regions. There are several metrics that are traditionally used in machine learning literature, namely Precision, which is basically a metric that quantifies the positive predictive value of the model, and Recall that measures the model’s sensitivity to detecting positive regions.
@@
To optimize computational efficiency, we initialized the encoder of the hybrid model with pre-trained weights derived from ImageNet[imagenet]. This transfer learning approach reduced convergence time and enhanced the model’s ability to extract domain-relevant features from limited annotated data, as previously demonstrated in medical image segmentation tasks [unet++]. The training process was conducted on GPUs provided by Kaggle, imposing constraints on computational resources. Nevertheless, through the use of pretraining and an empirically fine-tuned loss function, the currently employed architecture achieved a notably high level of accuracy while also being quite computationally efficient. The combined strategies ensure that the model is robust, efficient, and capable of performing effectively on local clinical datasets with inherent variability.
@@ 3 Results and Discussion
A comparative analysis against leading architectures is shown in Table 2. Aboelenein et al. (2020)[HTTU] proposed HTTU Net and reported a Dice Score of 0.865 and an IoU of 0.745 on the BraTS 2018 dataset, which comprises roughly 171 000 slices. Likewise, Li et al. (2018)[HDENSENET] developed H DenseUNet and achieved a Dice Score of 0.882 and an IoU of 0.792 on the LiTS 2017 dataset of 58 638 slices. Syazwany et al. (2021)[MMBiFPN] introduced MM BiFPN, obtaining a Dice Score of 0.8107 and an IoU of 0.7031 on the BraTS 2018 and 2020 datasets, which include about 171 000 and 221 400 slices respectively. By comparison, our final hybrid model integrates the Transformer Bottleneck, efficient attention[shen2021efficient] mechanisms, Squeeze and Excitation (SE)[hu2018squeeze] blocks, Convolutional Block Attention Module (CBAM)[woo2018cbam], and ResNeXt modules, achieving a Dice Score of 0.7636 and an IoU of 0.7357. It was trained on a local clinical dataset of only 6 080 slices. Considering the challenges of working with a small heterogeneous clinical dataset under limited computing resources, these results highlight the robustness and effectiveness of the proposed architecture.

### v2

@@ 1 INTRODUCTION
Improper classification and identification of brain tumors is a major cause of high mortality rate worldwide in afflicted patients. An estimated 4,407 deaths from brain and central nervous system cancers occurred in Pakistan in 2022, highlighting the critical need for improved diagnosis and treatment methods [1]. Magnetic resonance imaging (MRI) offers superior soft-tissue contrast, making it the modality of choice for identifying brain tumors and lesions [2]. Precise segmentation of these abnormalities is vital for both early detection and the formulation of appropriate treatment strategies. An ischemic stroke results from the obstruction of a cerebral blood vessel, which deprives brain tissue of essential oxygen and nutrients. Such an event represents a critical medical emergency that may lead to irreversible neurological damage or fatality[3]. Manual segmentation remains the clinical gold standard for delineating tumor boundaries, despite being time-consuming and subject to inter-observer variability[4]. Medical experts in local hospitals still use manual segmentation of tumor regions to mark areas and boundaries of damaged brain tissue.
@@
Manual segmentation has its own daunting challenges which impact the patient directly. It is a time-consuming process and often a difficult task left out for the oncologists. Unlike well-defined organs (eyes, brainstem), tumors are irregular and unpredictable. They have inconsistent shape, size, boundaries and often blend with surrounding tissues, making detection harder. Dynamic growth and irregular borders complicate precise segmentation, and the contouring is often subjective to the view of the oncologist. There is a need for automated methods for tumor segmentation that can offer a more efficient and automated approach to diagnosing and treating tumors[5]. Supervised deep learning methods like CNNs have contributed to image segmentation tasks in medical physics greatly over the years, especially reflecting in the BRATS[6] challenges annually. UNet [7] is one of the most used deep learning architecture in medical image segmentation. The UNET architecture consists of an encoder and decoder block linked with a bottleneck layer in a U shape (hence the name UNet). The purpose of the encoder block is to downsample the input image using convolutional and pooling layers to condense the information in it by contracting the spatial resolution and depth structures, passing it to a bottleneck layer that connects it to the decoder layer. The decoder layer then upsamples the condensed image to an output label mask using Transpose Convolutions.
@@
While pure CNNs excel at capturing local features, they struggle with long-range dependencies. A recent survey of self-supervised learning mechanisms for Vision Transformers demonstrates how SSL-pretrained ViTs can greatly reduce labeled data requirements while preserving downstream segmentation accuracy [8]. At the same time, an extensive review of Vision Transformer–based models for medical image segmentation reports that hybrid CNN–ViT architectures consistently outperform pure CNNs on MRI and CT benchmarks by better modeling global context [9].
@@
While models like UNET and its variants have shown remarkable success in the BRATS challenges, they face significant hurdles when applied to real-world clinical data, particularly in resource-limited settings. The reliance on curated datasets such as BRATS poses critical challenges for models aimed to be deployed locally in hospitals for radiation oncology treatment planning. Public datasets often fail to capture the heterogeneity of local patient populations, imaging protocols, and scanner variations. In addition, Models trained on BRATS or similar datasets may not generalize well to localized datasets, which are smaller in size and exhibit greater variability[10]. Data privacy concerns and regulations restrict the availability of medical data, thus limiting the prospect of such studies to test out models on vast amount of data[11].
@@ 2.1 Dataset and Data Preparation
The dataset used in this study was directly collected from a local hospital, offering a unique opportunity to evaluate our segmentation method on authentic clinical data. Unlike standardized benchmarks such as the BraTS dataset, which is employed in state‐of‐the‐art approaches like nnU-Net [13] and UNETR [14] or the MICCAI LiTS [15] dataset used in H-DenseUNet [16], our dataset reflects the natural variability encountered in routine clinical practice. The inherent differences in imaging protocols and patient demographics introduce challenges for segmentation models, while simultaneously enhancing the real-world relevance of our findings.
As summarized in Table 1, the dataset comprises MRI scans from six patients, with four cases acquired using T1-weighted sequences and two with T2-weighted sequences. Each patient contributed approximately 175 slices, yielding a total of 700 T1-weighted and 350 T2-weighted images. All images were acquired at a voxel resolution of acquired at a voxel resolution of . The raw data, stored as DICOM files, were initially converted into volumetric images using the open-source PyDICOM library [17]. Expert-provided segmentation masks were then transformed into binary label maps, and the volumetric data were sliced into 2D images, each paired with its corresponding mask. These processed images were subsequently organized into TensorFlow data pipelines to facilitate efficient model training and evaluation.
@@
Given the modest size of the original dataset (1̃,000 images), an extensive data augmentation strategy was implemented to enhance both data diversity and model robustness. As illustrated in Figure 2, several data augmentation techniques including flips, rotations, Gaussian blur, and contrast shifting are applied that expand the size of the dataset to 6,080 images. Such augmentation methods have been shown to improve the generalization and accuracy of deep learning models in medical imaging tasks [18].
@@ 2.2 Proposed Methodology
Vaswani et al [19] introduced the transformer architecture, which employs a self attention mechanism to process sequences concurrently, yielding high efficiency. This approach enables the network to attend to the most pertinent regions of an image by mapping a query against collections of keys and values to produce an output. A compatibility function between the query and each key computes weights for the associated values, which are then aggregated through a weighted sum to form the final output [19]. By assigning attention weights in this way, the model captures long range dependencies and attains a wider receptive field than convolutional neural networks, thereby integrating broader contextual information. Combining the merits of both transformer and convolutional designs, hybrid models have shown remarkable performance improvement over baseline models.
Hybrid architectures have emerged as a significant advancement in deep learning, particularly for medical image segmentation tasks[21]. By leveraging the strengths of multiple methodologies, these architectures aim to improve performance and address inherent limitations in conventional models. Traditional UNet architectures, renowned for their inductive bias and efficient handling of small datasets, often struggle to capture long-range dependencies[7]. Subsequently, transformer-based models excel in modeling global relationships but demand extensive annotated data, making them impractical for resource-constrained environments[22]. Hybrid models offer a balanced solution by integrating the complementary features of these approaches[14].
Several hybrid architectures have been proposed in the literature. The Hybrid UNet Transformer (HUT) [23] combines convolutional layers with transformers, introducing transformers at skip connections to enhance receptive fields. While effective in improving attention mechanisms, its reliance on large annotated datasets limits its applicability to resource-constrained settings. Similarly, the MaxViT-UNet framework [12] integrates MaxViT blocks into an encoder-decoder structure, achieving significant parameter efficiency and local-global feature modeling. However, the computational needs and resources required for such models make them less suitable for deployment in under-resourced environments. Another notable approach, the Two-Track UNet [24], utilizes dual convolutional tracks to address class imbalances, yet its reliance on specific kernel configurations reduces adaptability to heterogeneous datasets. Finally, H-DenseUNet [16] fuses 2D and 3D DenseUNets to handle intra- and inter-slice features, achieving strong performance on liver and tumor segmentation tasks. Despite its effectiveness, the model’s dependence on large-scale public datasets and computational resources creates challenges for real-world clinical deployment.
@@
Our proposed hybrid architecture (see Figure 4) builds upon the canonical U-Net framework to achieve precise segmentation of clinical images while addressing the limitations of resource-constrained training environments. In the Encoder, we utilize a ResNet-50[25] backbone from the [26] library, pre-trained on ImageNet[27]. This design choice is motivated not only by ResNet-50’s[25] proven ability to extract robust hierarchical features from input images but also by the practical necessity to reduce training time and computational cost on Kaggle GPUs, which impose strict limits on available resources. The encoder produces feature maps at four distinct stages: Stage 1 yields a tensor capturing low-level details (edges and textures), Stage 2 outputs a tensor representing medium-level features, Stage 3 produces a tensor encapsulating high-level information, and Stage 4 delivers a tensor that abstracts semantic content. To further enhance these representations, each encoder block is augmented with Squeeze-and-Excitation (SE)[28] and Convolutional Block Attention Modules (CBAM)[29]. The SE block, for example, recalibrates the input by first computing a channel descriptor via global average pooling, then passing it through two FC layers with a reduction ratio r (typically 16): and finally re-weighting the input by performing element-wise multiplication . This mechanism enhances the network’s sensitivity to clinically important features, such as subtle tumor boundaries.
@@
At the network’s core, the Transformer Bottleneck is employed to capture long-range dependencies that conventional convolutional layers cannot. Here, the deepest encoder output is a tensor, that is reshaped into a sequence of 1024 tokens (i.e., ) and processed through a series of four transformer blocks. Each block applies multi-head self-attention[19] (as described in the previously provided Equation 1 and a feedforward network with 1024 hidden units, along with dropout (0.1) and layer normalization. This design significantly broadens the receptive field and integrates global contextual information, which is essential for delineating diffuse and irregular tumor regions.
@@
In the Decoder, the network gradually upsamples the feature maps to restore the original resolution. At every upsampling stage, skip connections fuse high-resolution encoder features into the decoder output. Before fusion, the skip connection features are refined using CBAM to recalibrate channel and spatial responses. Moreover, an efficient attention[30] mechanism is applied within the decoder to dynamically emphasize the most informative features. In this module, the input tensor is first reduced in channel dimension using a convolution, followed by an additive branch and a ReLU activation; a subsequent convolution, followed by a sigmoid activation, yields attention coefficients that modulate the original feature map as , thereby filtering out irrelevant information. To further refine these features, each decoder block incorporates additional SE blocks and ResNeXt[31] modules. The ResNeXt block leverages grouped convolutions to capture multi-scale spatial patterns efficiently. It first reduces channel dimensions via a convolution, then processes spatial information using a depthwise convolution, and finally expands the channel dimension with another convolution. A parallel shortcut branch ensures that the block benefits from residual learning, and a final ReLU activation completes the operation.
@@
In the Decoder (right, shown in green), the network gradually upsamples the feature maps back into an image of original resolution. Skip connections, as indicated by the dashed arrows labeled “Attn Gate” in Figure Figure 4, fuse high-resolution encoder features with the decoder output. Prior to fusion, each skip connection is refined using CBAM, and then the decoder applies an efficient attention mechanism to selectively emphasize the most informative features. Furthermore, the decoder interleaves additional SE blocks and ResNeXt blocks; the latter leverages grouped convolutions[31] to capture multi-scale patterns efficiently, ensuring that both local details and global context are preserved. This comprehensive integration of modules enables our architecture to balance the data efficiency of convolutional networks[7] with the long-range image dependency modeling of transformers[21], thus achieving robust segmentation performance even on limited clinical datasets.
@@ 2.3 The Training of Hybrid Architecture
where is a hyperparameter empirically optimized to balance the contributions of the two losses. This hybrid formulation has demonstrated significant effectiveness in balancing global context with fine-grained segmentation accuracy in medical image segmentation tasks [32].
@@ 2.3.1 Evaluation Metrics
The segmentation performance is assessed based on some evaluation metrics, we employed several metrics commonly used in medical imaging studies [13, 24, 16]. These metrics include Dice Score, which is a measure of the degree of overlap between predicted and ground truth image, similar to Intersection over Union(IoU), which quantifies the ratio of the intersection to the union of predicted and actual regions. There are several metrics that are traditionally used in machine learning literature, namely Precision, which is basically a metric that quantifies the positive predictive value of the model, and Recall that measures the model’s sensitivity to detecting positive regions.
@@
To optimize computational efficiency, we initialized the encoder of the hybrid model with pre-trained weights derived from ImageNet[27]. This transfer learning approach reduced convergence time and enhanced the model’s ability to extract domain-relevant features from limited annotated data, as previously demonstrated in medical image segmentation tasks [33]. The training process was conducted on GPUs provided by Kaggle, imposing constraints on computational resources. Nevertheless, through the use of pretraining and an empirically fine-tuned loss function, the currently employed architecture achieved a notably high level of accuracy while also being quite computationally efficient. The combined strategies ensure that the model is robust, efficient, and capable of performing effectively on local clinical datasets with inherent variability.
@@ 3 Results and Discussion
A comparative analysis against leading architectures is shown in Table 2. Aboelenein et al. (2020)[24] proposed HTTU Net and reported a Dice Score of 0.865 and an IoU of 0.745 on the BraTS 2018 dataset, which comprises roughly 171 000 slices. Likewise, Li et al. (2018)[16] developed H DenseUNet and achieved a Dice Score of 0.882 and an IoU of 0.792 on the LiTS 2017 dataset of 58 638 slices. Syazwany et al. (2021)[34] introduced MM BiFPN, obtaining a Dice Score of 0.8107 and an IoU of 0.7031 on the BraTS 2018 and 2020 datasets, which include about 171 000 and 221 400 slices respectively. By comparison, our final hybrid model integrates the Transformer Bottleneck, efficient attention[30] mechanisms, Squeeze and Excitation (SE)[28] blocks, Convolutional Block Attention Module (CBAM)[29], and ResNeXt modules, achieving a Dice Score of 0.7636 and an IoU of 0.7357. It was trained on a local clinical dataset of only 6 080 slices. Considering the challenges of working with a small heterogeneous clinical dataset under limited computing resources, these results highlight the robustness and effectiveness of the proposed architecture.